# Beyond Conventional Transformers: The Medical X-ray Attention (MXA) Block for Improved Multi-Label Diagnosis Using Knowledge Distillation

### v1

@@ 3.3 Training and Evaluation
and the weights of EfficientViTs are updated via standard backpropagation.
During validation, we compute the same loss. Additionally, to track a simple measure of correctness for each label [31], we compute an element-wise comparison of $\sigma(O)$ with $Y$, thresholding at $0.5$. The resulting “accuracy” reflects the fraction of individual pathology labels predicted correctly across the batch.
3.4 Knowledge Distillation with a Multi-label Teacher
To further improve performance, especially for challenging or subtle pathologies, we apply knowledge distillation (KD) in a multi-label context. Our DistillationLoss class uses a teacher model (a pre-trained DenseNet-121 on CheXpert with 14 outputs) and compares the student’s predictions to the teacher’s logits transformed by a sigmoid. Formally, let $O^{s}$ by the student’s output and $O^{t}$ be the teacher’s output. TEH teacher probabilities are
$p^{t}=\sigma(O^{t})$
and we compute a distillation loss that is either:
  1. 1.
    Soft Distillation: Using the teacher probabilities directly in a BCE-like loss against the student logits:
    $\mathscr{L}^{\text{soft}}_{\text{distill}}=\text{BCEWithLogitsLoss}(O^{s},p^{t})$
  2. 2.
    Hard Distillation: Binarize the teacher probabilities at $0.5$ to form pseudo-labels $h^{t}=\mathbb{I}\bigl{[}p_{t}>0.5\bigr{]}$, then compute a BCE loss between $O^{s}$ and $h^{t}$.
In both cases, the total loss is a weighted combination of the base BCEWithLogitsLoss and $\mathscr{L}_{\text{distill}}$. Formally, if $\alpha$ is the distillation weight,
$\mathscr{L}_{\text{total}}=(1-\alpha)\mathscr{L}_{\text{BCEWithLogits}}(O^{s},Y)+\alpha\mathscr{L}_{\text{distill}}(O^{s},O^{t})$
Through this multi-label KD strategy, the student model more effectively learns to approximate the teacher’s predictions, leading to improved accuracy.
4 Medical X-ray Attention (MXA) Block
Figure 1: Integration of the Medical X-ray Attention (MXA) Block Architecture.
4.1 Dynamic Region-of-Interest (ROI) Pooling
Multi-label X-ray classification often includes multiple abnormal regions and purely global feature pooling might dilute small or subtle findings. To identify the most relevant regions for targeted analysis, consequently reducing computational overhead, we propose a Dynamic ROI pooling approach. By dynamically pooling region-of-interest (ROI) features, the model subsectionasizes critical areas while ignoring extraneous regions.
Given an input feature map $F\in\mathbb{R}^{BxCxHxW}$, where $B$ is the batch size, $C$ is the number of channels, $H,W$ are the height and width of the feature map, our ROI pooling mechanism selects critical regions using a region predictor, formally defined as:
$\text{ROI}_{i}=[x_{1},y_{1},x_{2},y_{2}],\ \text{where}\ i=1,2,\ldots,B$
Here, $[x_{1},y_{1}]$ and $[x_{2},y_{2}]$ are normalized coordinates of the top-left and bottom-right corners of the region of interest (ROI) for the $i$-th image. These coordinates are generated by a lightweight convolutional network conditioned on the feature map. The bounding-box region predictor is trained jointly with the classification objective in an end-to-end fashion: during the forward pass, each predicted ROI is used to crop and rescale the input feature map, and the resulting ROI features flow into the rest of the network for classification. The gradients from the multi-label classification loss then back-propagate through the ROI predictor as well, optimizing both the bounding boxes and the classification parameters simultaneously. Consequently, the network learns ROIs that best localize relevant abnormalities for improved multi-pathology detection.
After cropping the predicted ROI from the feature map, we resize it to match the original feature map size via bilinear interpolation:
$F^{(i)}_{\text{pooled}}=\text{Resize}(F_{\text{ROI}_{i}},(H,W))$
where $F^{(i)}_{\text{pooled}}$ is the pooled ROI feature map, and Resize()($\cdot$)( ⋅ ) denotes the bilinear interpolation function. By pooling only the relevant region to full resolution, we reduce the overhead that would otherwise come from processing the entire X-ray at high granularity [32]. This ensures compatibility with subsequent operations while focusing on the most critical regions. We use a kernel size of 3 in certain layers (e.g. local window attention or ROI-based convolutional components) as a practical middle ground for ROI between capturing sufficient local context and avoiding over-fitting in chext X-ray data. Prior work in medical image analysis (e.g., [6]) has similarly found that intermediate kernel sizes strike a good balance between detail preservation and robust feature extraction.
4.2 CBAM-Style Attention
To further refine the representation and enhance localized feature learning, we incorporate a Convolutional Block Attention Module (CBAM) [18] into the MXA block. CBAM sequentially applies channel attention and spatial attention to the pooled feature maps, enabling the network to focus on the most informative channels and spatial regions. This combined attention is especially helps the network prioritize more pathologically relevant features, crucial in medical images where anomalies can be small or faint. Spatial in particular helps improve the overall global context.
Channel attention determines which feature channels contribute the most to the classification task. This is achieved by globally aggregating the feature map along the spatial dimensions using global average pooling (GAP) and global max pooling (GMP), followed by a shared multi-layer perceptron (MLP) [33]. For a given pooled feature map $F_{pooled}\in\mathbb{R}^{C\times H\times W}$, channel attention is computed as:
$M_{C}=\sigma(W_{2}\delta(W_{1}\text{GAP}(F_{\text{pooled}}))+W_{2}\delta(W_{1}\text{GMP}(F_{\text{pooled}})))$
where $W_{1}\in\mathbb{R}^{C\times\frac{C}{r}}$ and $W_{2}\in\mathbb{R}^{\frac{C}{r}\times C}$ are learnable weight matrices, $\delta$ is the ReLU activation function [34], $\sigma$ is the sigmoid function [35], and $r$ is the reduction ratio, controlling the dimensionality reduction in the MLP. The resulting attention map $M_{C}\in\mathbb{R}^{C}$ is applied to teh feature map by element-wise multiplication
$F_{\text{channel-attention}}=M_{C}\times F_{\text{pooled}}$
Next, spatial attention identifies the most critical regions within each channel-refined feature map [18]. Using the channel-refined feature map $F_{\text{channel-attention}}$, it pools along the channel dimension using max pooling and average pooling to generate two spatial descriptors:
$F_{\text{spatial}}=[\text{MaxPool}(F_{\text{channel-attention}}),\text{AvgPool}(F_{\text{channel-attention}})]$
where $F_{\text{spatial}}\in\mathbb{R}^{2\times H\times W}$. These pooled descriptors are concatenated and passed through a convolutional layer to produce the spatial attention map $M_{S}$:
$M_{S}=\sigma(\text{Conv2D}(F_{\text{spatial}}))$
where $M_{S}\in\mathbb{R}^{H\times W}.$ The spatial attention map is applied to the feature map by element-wise multiplication:
$F_{\text{spatial-attention}}=M_{S}\times F_{\text{channel-attention}}$
$F_{\text{CBAM-attention}}=F_{\text{spatial-attention}}$
The combined effect of channel and spatial attention refines the feature representation by prioritizing both the most informative channels and regions. Incorporating CBAM into the MXA block provides a richer and more adaptive representation of X-ray feature maps.
4.3 Parallel Integration with MHSA
To further enhance feature representation and ensure that the model effectively utilizes both global dependencies (captured by MHSA [1]) and localized features (captured by MXA) are effectively utilized, we propose integrating the MXA block in parallel with the MHSA layers in EfficientViTs.
$F_{\text{output}}=F_{\text{MHSA}}+F_{\text{MXA}}$
where $F_{\text{MHSA}}$ represents the output from the MHSA layer, and $F_{\text{MXA}}$ is the refined feature map from the MXA block. Refer to Fig. 1 for an illustration of the MXA block and its integration into the overall transformer framework.
(a) pt1
(b) pt2
(c) pt3
(d) pt4
(e) pt5
(f) pt6
(g) pt7
(h) pt8
Figure 2: Demonstration of the MXA Block. Each subfigure (pt1–pt8) illustrates the region of interest dynamically pooled by the MXA block after an initial MXA inference run for individual chest X-rays. All patients presented exhibit a pneumonia pathology as a shared characteristic. The red bounding boxes denote the predicted ROI coordinates identified by the MXA block. Associated metadata for additional pathologies (e.g., Pneumothorax, Cardiomegaly) are detailed in Table 1. In accordance with ethical standards and the CheXpert Terms of Service, non-essential metadata has been omitted. This visualization underscores MXA’s capability to highlight critical local features while simultaneously maintaining global contextual awareness.
Table 1: Metadata for the 8 MXA Test Figures
Patient NF ECM CM LO LL ED CON PNA ATL PTX PE PO FX SD
pt1 0 1 1 1 0 0 1 1 1 0 1 0 0 0
pt2 0 1 1 1 0 1 1 1 1 0 1 0 0 0
pt3 0 0 0 1 0 0 1 1 0 0 0 0 0 1
pt4 0 0 0 1 0 0 1 1 1 0 1 0 0 0
pt5 0 0 0 1 0 0 1 1 1 0 1 1 0 1
pt6 0 1 1 1 0 0 1 1 1 0 1 0 0 0
pt7 0 1 1 1 0 0 1 1 1 0 1 0 0 1
pt8 0 0 0 1 0 1 1 1 1 0 1 0 0 0
5 Knowledge Distillation
Knowledge Distillation (KD) is a widely used technique for transferring knowledge from a large, well-trained teacher model to a more compact student model [19]. In our work, KD is employed within the EfficientViT framework to enhance the recognition of subtle and co-occurring pathologies in multi-label chest X-ray classification. In the following subsections, we detail the teacher model, the transformation layer and label mapping that adapts the teacher’s outputs, our distillation strategy, and the dynamic label weighting mechanism used to refine the supervisory signal.
We adopt a DenseNet-121 [26, 9] teacher model from TorchXRayVision [^1] https://github.com/mlmed/torchxrayvision (specifically the “densenet121-res224-chex” variant [36]), pre-trained on the CheXpert dataset [16] with 18 output nodes. In contrast, our student model is designed for 14 pathology labels. The discrepancy in label formats arises from the teacher model including additional categories (e.g., Support Devices and Pleural Other) that may be uninitialized or perform suboptimally. This motivates the need for a transformation layer to realign the teacher outputs with the student’s label space.
5.1 Transformation Layer and Label Mapping
To reconcile the teacher’s 18-label output with the 14-label format required by our student model, we introduce a transformation layer. This layer performs the following operations:
  • It rearranges the teacher logits to match the ordering expected by the student. The module uses a predefined mapping (e.g., student index 1 maps to teacher index 17) to correctly align the outputs.
  • The adapter computes the "No Finding" score as the product of the complements of the sigmoid-transformed teacher logits:
    No $\text{No Finding}=\prod_{i=1}^{N}(1-\sigma(O^{t}_{i}))$
    ensuring that this aggregated score accurately reflects the absence of pathologies.
  • By explicitly setting the logit corresponding to non-existent labels in the student model to 0, the adapter effectively ignores these unreliable outputs during the loss computation.
The functionality is implemented in our TeacherOutputAdapter module, which maps and filters the teacher logits accordingly.
5.2 Soft Distillation and Dynamic Weighting
In our KD framework, the student model’s output $O^{s}$ is trained to mimic the teacher’s adapted outputs. The teacher logits $O^{t}$ are first transformed and then passed through a sigmoid function to obtain probabilities:
$p^{t}=\sigma(O^{t})$
In our experiments, soft distillation is employed since it better captures the nuanced confidence levels of the teacher’s outputs. To further refine the distillation process, we incorporate dynamic label weighting based on the teacher’s confidence. Let $p^{t}_{ij}$ denote the teacher’s predicted probability for label $j$ on the $i$-th sample in a batch of size $B$. The dynamic weight $w_{j}$ for label $j$ is computed as:
$w_{j}=1-\frac{1}{B}\sum_{i=1}^{B}p^{t}_{ij}$
This formulation assigns lower weights to labels for which the teacher exhibits high confidence and higher weights to more uncertain labels, thereby emphasizing learning where the teacher is less certain. These weights are integrated into the distillation loss to focus the training on challenging or ambiguous pathologies.
5.3 Overall Loss and Training Dynamics
The final loss function is formulated as a weighted sum of the standard BCEWithLogitsLoss on the ground-truth labels $Y$ and the distillation loss:
$\mathscr{L}_{\text{total}}=(1-\alpha)\,\mathscr{L}_{\text{BCEWithLogits}}(O^{s},Y)+\alpha\,\mathscr{L}^{\text{soft}}_{\text{distill}}(O^{s},\,p^{t})$
where $\alpha$ is the distillation weight. While single-label KD often uses Kullback–Leibler divergence on softmax outputs [14], our setting requires BCE due to multiple independent pathologies. To accommodate this, we introduce a customized DistillationLoss class (refer to Section 3.4). The integration of the transformation layer and dynamic label weighting ensures that unreliable teacher outputs are effectively ignored while reinforcing learning from more dependable signals. Gradients propagate through the entire EfficientViT student network during training, while the teacher model remains frozen. This strategy allows the student to effectively learn the complex multi-label patterns inherent in chest X-ray data, resulting in improved diagnostic performance.
@@ 6 Experiments
We evaluate our multi-label EfficientViT models for chest X-ray classification, comparing a naive multi-label EfficientViT baseline against our proposed method that integrates knowledge distillation (KD) and the MXA block. In both cases, models are trained and validated on the CheXpert dataset [16], and performance is reported via standard multi-label metrics (average accuracy and F1-score) at a threshold of 0.5.
6.1 Multi-label EfficientViT Network Architectures
Table 1 outlines the principal variants of our multi-label EfficientViT architectures. Each model family member differs in its stage-wise width ($C_{i}$), depth ($L_{i}$), and number of heads ($H_{i}$). Specifically, we break the network into three stages, and each stage’s dimensions are adjusted to control the overall capacity and computational cost.
Table 2: Multi-label EfficientViT Architecture Variants
Model $\{C_{1},C_{2},C_{3}\}$$\{L_{1},L_{2},L_{3}\}$$\{H_{1},H_{2},H_{3}\}$
EfficientViT-MultiLabel-M0 {64, 128, 192} {1, 2, 3} {4, 4, 4}
EfficientViT-MultiLabel-M1 {128, 144, 192} {1, 2, 3} {2, 3, 3}
EfficientViT-MultiLabel-M2 {128, 192, 224} {1, 2, 3} {4, 3, 2}
EfficientViT-MultiLabel-M3 {128, 240, 320} {1, 2, 3} {4, 3, 4}
EfficientViT-MultiLabel-M4 {128, 256, 384} {1, 2, 3} {4, 4, 4}
EfficientViT-MultiLabel-M5 {192, 288, 384} {1, 3, 4} {3, 3, 4}
@@ 6.2 Implementation Details
All models are implemented in PyTorch 2.5.1 [37] using the Timm 0.5.4 library [38]. Training is performed from scratch for 50 epochs on a single NVIDIA H100 GPU, employing the AdamW optimizer [39] with a weight decay of 0.025 and the ReLU activation function within all all EfficientViT blocks.. We use the preset train-validation split provided by CheXpert, ensuring consistency with prior benchmarks. Chest X-ray images are resized to 224 × 224, and we use a 16 × 16 patch size for tokenization. The LocalWindowAttention module within each EfficientViT block uses a $7\times 7$ local window, balancing local context capture with computational efficiency.
We train with a batch size of 512 and set an initial learning rate of $1\times 10^{-3}$, following a cosine learning rate scheduler with a minimum learning rate of $1\times 10^{-5}$. Gradient clipping is applied with a maximum norm of 0.02 using adaptive gradient clipping (AGC). The learning rate warmup lasts for 5 epochs, and the cosine decay rate is 0.1 over 30 decay epochs, followed by 10 cooldown epochs. EMA (Exponential Moving Average) is enabled with a decay factor of 0.99996.
For knowledge distillation (KD), we employ DenseNet-121 [26] (“densenet121-res224-chex” variant [36]) pretrained on CheXpert (with the original 14-label output). We remove its final classification layer, fine-tune it for multi-label classification, and keep it frozen during student model training. Soft distillation is applied with $\alpha=0.5$ and a temperature $(\tau)$ of 1.0. The final classification head in each EfficientViT variant produces 14 outputs, corresponding to the same set of pathologies. We use bicubic interpolation for image resizing during training and apply RandAugment (rand-m5-mstd0.2-inc2) for data augmentation. Stochastic depth, dropout, mixup, and cutmix are disabled to ensure stable training. The full architecture details of our model family are presented in Table 1.
@@ 6.3 Training Protocol
Baseline Method: Multi-label Efficient Vision Image Transformer
  1. 1.
    Architecture: We employ EfficientViT blocks (with LocalWindowAttention) in three stages, each with its own depth and embedding dimension (see Tab 1.).
  2. 2.
    Loss: BCEWithLogitsLoss for multi-label classification, summing over the 14 pathologies.
  3. 3.
    Learning Rate: We use a cosine learning rate scheduler starting at $31\times 10^{-}$ 3.
  4. 4.
    No Distillation, No MXA: The model is purely a multi-label extension of EfficientViT, using a patch embedding and standard cascaded group attention without the new proposed modules.
Proposed Method: Multi-label Efficient Vision Image Transformer with the Medical X-ray Attention Block and Knowledge Distillation
  1. 1.
    MXA Block: In each of the three stages, we add an MXA module in parallel to the MHSA. MXA learns dynamic ROIs and integrates additional channel/spatial attention (via CBAM) to subsectionasize critical regions.
  2. 2.
    Knowledge Distillation
    • Teacher: DenseNet-121 pretrained on CheXpert.
    • Student: Our multi-label EfficientViT with the MXA block.
    • Loss: Weighted sum of BCEWithLogitsLoss (ground-truth) and distillation loss $\mathscr{L}_{\text{distill}}$ from teacher probabilities.
  3. 3.
    ROI Predictor: Trained end-to-end using bounding-box predictions that highlight suspicious or abnormal regions, rescaled to full size via bilinear interpolation for further attention processing.
  4. 4.
    Training Scheme: We use the same optimizer and learning rate schedule as the naïve approach. The teacher’s weights are kept frozen, and only the student updates its parameters.
@@ 7 Results
We evaluate and report our main findings on the CheXpert dataset, comparing the proposed EfficientViT with the Medical X-ray Attention Block and Knowledge Distillation approach against a baseline multi-label EfficientViT. We further present an ablation study in which different components (the Medical X-ray Attention block and knowledge distillation) are incrementally introduced. Finally, we analyze the per-pathology performance to highlight the strengths and weaknesses of our approach for specific conditions.
@@ 7.1 Overall Performance
Table 3 compares the baseline multi-label EfficientViT model with our proposed method, which integrates the MXA block and knowledge distillation across 50 epochs on validation. The baseline achieves an ROC AUC of 0.66, whereas our proposed method attains an ROC AUC of 0.85. Notably, this improvement of 0.19 in absolute terms (66% to 85%) or relative improvement of roughly 29%, corresponds to a large gain relative to the random-guessing baseline of 50% AUC. We also observe consistent improvements in the average accuracy, F1-score, and cross-entropy loss, demonstrating that our model captures both robust global signals (via the transformer backbone) and subtle local anomalies (through MXA).
Table 3: Overall Performance Comparison
Model Accuracy (%) Loss F1-Score ROC AUC
Baseline (EfficientViT) 84.0 0.679 0.476 0.661
Proposed (EfficientViT + MXA + KD) 84.4 0.406 0.599 0.8529
Figure 3: Boxplot comparison of the AUC per epoch for three independent test runs and their aggregate performance. Each subplot compares the baseline model with the proposed model. The proposed model consistently outperforms the baseline across all runs, demonstrating robustness and generalization improvements.
Figure 4: Bar plot of the average AUC over three independent runs for the baseline and proposed models. Error bars represent a 95% confidence interval (approximated as ±$\pm 2\sigma$). The proposed model shows a significantly higher and more stable performance across runs.

### v2

@@ 3.3 Knowledge Distillation
where $B$ is the batch size and $p^{\mathrm{t}}_{ij}$ the teacher probability for label $j$ on sample $i$. These weights down-scale loss contributions from high-confidence teacher predictions $(p^{\mathrm{t}}_{ij}\!\approx\!1)$ and amplify ambiguous cases; refer to Appendix E for the rigorous soft distillation term.
@@ 4 Experiments
We evaluate our multi-label M5 EfficientViT for CXR classification, comparing a naive multi-label EfficientViT baseline against our proposed method that integrates KD and the MXA block. In both cases, models are trained and validated on the CheXpert dataset [6], and performance is reported via standard multi-label metrics (average accuracy, F1-score, and micro averaged AUC) at a positive label threshold of 0.5.
@@ 4.1 Implementation Details
The EfficientViT-MultiLabel-M5 is implemented in PyTorch 2.5.1 [27] using the timm 0.5.4 library [28]. Training is performed from scratch for 50 epochs on a single NVIDIA H100 GPU, employing the AdamW optimizer [29] with a weight decay of 0.025 and the ReLU activation function within all EfficientViT blocks. We use the preset train–validation split provided by CheXpert, ensuring consistency with prior benchmarks. Chest X-ray images are resized to $224\times 224$, and a $16\times 16$ patch size is used for tokenization. The LocalWindowAttention module within each M5 EfficientViT block employs a $7\times 7$ local window, balancing local context capture with computational efficiency.
We train with a batch size of 512 and set an initial learning rate of $1\times 10^{-3}$, following a cosine scheduler with a minimum of $1\times 10^{-5}$. Gradient clipping is applied with a maximum norm of 0.02 using adaptive gradient clipping (AGC). The learning-rate warm-up lasts for 5 epochs; cosine decay (rate 0.1) spans 30 decay epochs and is followed by 10 cool-down epochs. EMA (exponential moving average) is enabled with a decay factor of 0.99996.
For KD we employ DenseNet-121 [17] (“densenet121-res224-chex” variant [26]) pretrained on CheXpert. Its final classification layer is removed, the network is fine-tuned for multi-label classification, and kept frozen during student training. Soft distillation is applied with $\alpha=0.5$ and temperature $\tau=1.0$. Our EfficientViT-MultiLabel-M5 backbone (full spec in Appendix A, Table A.1) outputs 14 logits (one per pathology). Bi-cubic interpolation is used for resizing, with data augmentation via RandAugment (“rand-m5-mstd0.2-inc2”). Stochastic depth, dropout, mixup, and cutmix are disabled for stable training. EfficientViT-MultiLabel-M5 uses stage widths $\{192,288,384\}$, depths $\{1,3,4\}$, and heads $\{3,3,4\}$.
@@ 4.2 Training Protocol
Baseline (EfficientViT).
We train the three-stage M5 EfficientViT backbone with the AdamW optimizer and a cosine learning-rate schedule initialized at $1\times 10^{-3}$. The objective is the sum of $\operatorname{BCEWithLogitsLoss}$ over the 14 thoracic findings. No KD) or MXA modules are included, so the network reduces to a straightforward multi-label EfficientViT with patch embeddings and cascaded group attention.
Proposed (EfficientViT+MXA+KD).
Each stage now contains a parallel MXA block that learns dynamic regions of interest and applies CBAM-style channel–spatial attention to emphasize critical areas. A DenseNet-121 pretrained on CheXpert (frozen) acts as teacher, while the student minimizes a weighted sum of the ground-truth $\operatorname{BCEWithLogitsLoss}$ and the KD term $\mathcal{L}_{\text{distill}}$ computed from teacher probabilities. An ROI predictor is trained end-to-end, producing bounding boxes that are bilinearly up sampled before further attention processing. The optimizer and learning-rate schedule match the baseline, with updates applied only to student parameters.
@@ 5 Results
We evaluate and report our main findings on the CheXpert dataset, comparing the proposed M5 EfficientViT with the MXA Block and KD approach against a baseline multi-label M5 EfficientViT. We further present an ablation study in which different components (the MXA block and KD) are incrementally introduced. Finally, we analyze the per-pathology performance to highlight the strengths and weaknesses of our approach for specific conditions.
Table 1: Overall Performance Comparison
Model Accuracy (%) Loss F1-Score AUC
Baseline (M5 EfficientViT) 84.0 0.679 0.476 0.661
Proposed (M5 EfficientViT + MXA + KD) 84.4 0.406 0.599 0.8529
@@ 5.1 Overall Performance
Table 1 compares the baseline multi-label M5 EfficientViT model with our proposed method, which integrates the MXA block and KD across 50 epochs on validation. The baseline achieves an ROC AUC of $0.66$, whereas our proposed method attains an ROC AUC of $0.85$. Notably, this improvement of $0.19$ in absolute terms ($66\%\to 85\%$), a relative gain of roughly $29\%$, corresponds to a large margin over the random-guessing baseline of $50\%$ AUC. We also observe consistent improvements in the average accuracy, F1-score, and cross-entropy loss, demonstrating that our model captures both robust global signals (via the transformer backbone) and subtle local anomalies (through MXA). Figure 2 presents AUC (mean ±$\pm 1\sigma$) per epoch over three runs. The proposed model climbs above $0.8$ by epoch 10, compared to $0.65$ for the baseline, demonstrating faster early learning. Both curves plateau by epoch 30 with $\sigma<1.2\times 10^{-6}$, indicating stable convergence. A consistent $\approx 0.19$ micro-AUC gap persists throughout training, evidencing robust, enduring gains from MXA and KD.
Figure 2: Training AUC over 30 epochs for baseline and proposed models, showing stable convergence and a consistent performance margin. The improved model converges faster and sustains 0.19 higher AUC than the naive baseline throughout training, evidencing durable gains.